Circular No. 5720

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993G IN NGC 3690 A. V. Filippenko, T. Matheson, B. Leibundgut and L. C. Ho, Univer-

sity of California at Berkeley, report that CCD spectra (range 300-1000 nm) obtained on Mar. 13 with the Shane 3-m reflector at the Lick Observatory reveal that SN 1993G (cf. IAUC 5718, 5719) is of type II, within a few weeks of maximum brightness. The prominent Hα emission line, which peaks near 657 nm, is broad (FWHM ∼ 10 000 km/s) and asymmetric, but the corresponding absorption is weak. Hβ, Hγ and Hδ, by contrast, exhibit deep absorption. Narrow Hα emission from a faint underlying H II region appears at 663 nm (cz = 3070 km/s).

B. Schmidt and R. Kirshner, Center for Astrophysics, report that observations on Mar. 15 by S. Raychaudhury and G. Bernstein with the Multiple-Mirror Telescope confirm the supernova to be of type II. Hα, Hβ and He I (587.6 nm) show P-Cyg profiles, and absorption lines of Fe II (492.4, 501.8 and 516.9 nm) give a photospheric expansion of ∼ 600 km/s.

γ-RAY BURST A. Goldwurm, M. C. Schmitz-Fraysse, E. Churazov, M. Gilfanov, on

behalf of the SIGMA/Granat team (Service d’Astrophysique du CEA, Saclay; Centre d’Etude Spatiale des Rayonnements, Toulouse; Space Research Institute, Moscow), report: “The SIGMA telescope on board the Granat Observatory detected a γ-ray burst on Mar. 10.30576 UT. The event lasted ∼ 20 s and was detected in the 35-400-keV energy range through an interruption of the passive shield acting as a collimator. The localization process is based on a maximum-likelihood method yielding a signal-to-noise ratio of 16. The 1σ error box of the burst is approximately diamond-shaped, the positions of the corners being α = 19^{h}16^{m}01^{s}, δ = −49^{o}24′; 19^{h}04^{m}37^{s}, −52^{o}23′; 18^{h}49^{m}59^{s}, −54^{o}01′; 18^{h}58^{m}19^{s}, −52^{o}07′ (equinox 1950.0).”

NOVA CYGNI 1992 Further visual magnitude estimates: Feb. 8.98 UT, 10.7 (B. H. Granslo,

Fjellhamar, Norway); 13.97, 10.7 (H. Dahle, Oslo, Norway); 19.04, 10.8 (Granslo); 23.75, 10.5 (M. Tombelli, Montelupo, Italy); Mar. 8.48, 11.4 (W. G. Dillon, Missouri City, TX); 13.16, 11.2 (A. Boattini, Florence, Italy).

1993 March 16 Brian G. Marsden